OpenStudy (anonymous):

square root of -12 divided by 2?

OpenStudy (agreene):

\[\sqrt{\frac{-12}2}=i\sqrt{6}\] \[\frac{\sqrt{-12}} 2=i\sqrt3\] not sure which one you are asking.

OpenStudy (saifoo.khan):

i misread. sorry.